OpenStudy (lynda.boyce):

What is the rate of change of the revenue when there are 1000 downloads. So far I have R'(1000) = (130 - .019(1000))e^(-.000250(1000)). How do I calculate to get the answer?

OpenStudy (anonymous):

plug 1000 in for x

OpenStudy (welshfella):

you'll need a scientific calculator to do that one.

OpenStudy (lynda.boyce):

I have one, but how do I type it in?

OpenStudy (welshfella):

exactly as you've written it

OpenStudy (lynda.boyce):

I don't need to put in 2.718282 anywhere?

OpenStudy (welshfella):

no - just the e^x key

OpenStudy (lynda.boyce):

would I type in 130-.01981000e^-0.000150*1000

OpenStudy (lynda.boyce):

the answer I got was 130.00

OpenStudy (welshfella):

well you can simplify it by doing some of it manually first 130 - .019*1000 = 130 - 19 = 111 and the exponent of e = 0.000250*1000 = 0.25 so it boils down to 111* e^(-0.25) try that one

OpenStudy (welshfella):

130 is incorrect

OpenStudy (welshfella):

- you didnt put the parentheses in

OpenStudy (lynda.boyce):

27.75

OpenStudy (welshfella):

try again i got 86.45

OpenStudy (welshfella):

dont forget the negative and the parentheses

OpenStudy (lynda.boyce):

ok, I used a different calculator and got 86.4469

OpenStudy (welshfella):

yes thats right